Diagonal brace. The width of a rectangular gate is 2 meters (m) larger than its height. The diagonal brace measures sqrt of 6 meters.
Length x + 2
Height x
Diagonal Brace sqrt of 6.

Answer by dabanfield(803)

So you have a right triangle with sides x and x+2 and hypotenuse sqrt(6). By the Pythagorean Theorem:
x^2 + (x+2)^2 = (sqrt(6))^2
x^2 + x^2 + 4x + 4 = 6
2x^2 + 4x - 2 = 0
Divide both sides by 2:
x^2 + 2x - 1 = 0
Use the quadratic formula to find x and calculate x+2.